likeabossssssss (likeabossssssss):

Which ratio correctly compares 15 m to 45 cm? A. 10 : 3 B. 1 : 300 C. 100 : 3 D. 1 : 3 I think it is b

likeabossssssss (likeabossssssss):

or a

OpenStudy (damonte212):

15m×100cmm=1500cm 1500cm:45cm=1500cm15:45cm15=100:3 C.

likeabossssssss (likeabossssssss):

ok thx

likeabossssssss (likeabossssssss):

@tkhunny is it right?

OpenStudy (tkhunny):

Why not convert carefully and see? \(\dfrac{15\;m}{45\;cm} = \dfrac{15\;m\cdot\dfrac{100\;cm}{1\;m}}{45\;cm} = \dfrac{1500\;cm}{45\;cm} = \dfrac{1500}{45} = \dfrac{300}{9} = \dfrac{100}{3}\) Good work.